# Efficient Large Language Models: A Survey

### v2

@@ 3.2.2. Prompt Compression
Prompt compression (Figure 19(a)) accelerates the processing of LLM inputs through either condensing lengthy prompt inputs or learning compact prompt representations. Mu et al. (2023) propose to train LLMs to distill prompts into a more concise set of tokens, referred to as gist tokens. These gist tokens encapsulate the knowledge of the original prompt and can be stored for future use. In doing so, it is able to compress prompts by up to 26 times, leading to a reduction in floating-point operations per second (FLOPs) by up to 40%. Chevalier et al. (2023) propose AutoCompressors to condense long textual contexts into compact vectors, known as summary vectors, which can then be used as soft prompts for the language model. These summary vectors extend the model’s context window, allowing it to handle longer documents with much less computational cost. Jung and Kim (2023) propose Prompt Compression with Reinforcement Learning (PCRL) that employs a policy network to directly edit prompts, aiming to reduce token count while preserving performance. It achieves an average reduction of 24.6% in token count across various instruction prompts. Ge et al. (2023) propose In-context Autoencoder (ICAE), which consists of a learnable encoder and a fixed decoder. The encoder compresses a long context into a limited number of memory slots, which the target language model can then condition on. With such design, ICAE is able to obtain 4x context compression. Nugget 2D (Qin et al., 2023a) represents the historical context as compact “nuggets” that are trained to enable reconstruction. Furthermore, it has the flexibility to be initialized using readily available models like LLaMA. Lastly, LongLLMLingua (Jiang et al., 2023c) introduces a prompt compression technique containing question-aware coarse-to-fine compression, document reordering, dynamic compression ratios, and post-compression sub-sequence recovery to enhance LLMs’ key information perception.
3.2.3. Prompt Generation
@@
Prompt generation (Figure 19(b)) enhances the efficiency by automatically creating effective prompts that guide the model in generating specific and relevant responses instead of manual annotated data. AutoPrompt (Shin et al., 2020) proposes an automated method to generate prompts for a diverse set of tasks based on a gradient-guided search. It underscores the significance of human-written text in refining the quality and authenticity of data, emphasizing its pivotal role in optimizing LLM performance. TempLM (Zhang et al., 2022c) proposes to combine generative and template-based methodologies to distill LLMs into template-based generators, offering a harmonized solution for data-to-text tasks. PromptGen (Zhang et al., 2022a) is the first work considering dynamic prompt generation for knowledge probing, based on a pre-trained LLMs. It can automatically generate prompts conditional on the input sentence and outperforms AutoPrompt on the on the LAMA benchmark.
| Framework | Training | Fine-Tuning | Inference | Features |
DeepSpeed |
@@
Data Parallelism, Model Parallelism, Pipeline Parallelism, Prompt Batching, Quantisation, Kernel Optimizations, Compression, Mixture of Experts. |
Megatron |
@@
Data Parallelism, Model Parallelism, Pipeline Parallelism, Prompt Batching, Automatic Mixed precision, Selective activation Recomputation |
Alpa |
@@
Data Parallelism, Model Parallelism, Pipeline Parallelism, Operator Parallelism, Automated Model-Parallel Training, Prompt Batching |
Colossal AI |
@@
Data Parallelism, Model Parallelism, Pipeline Parallelism, Mixed Precision Training, Gradient accumulation, heterogeneous Distributed Training, Prompt Batching, Quantization |
FairScale |
@@
Data Parallelism, Model Parallelism, Pipeline Parallelism, Activation Checkpointing, Model Offloading, Model scaling, Adascale Optimization |
Pax |
@@
Data Parallelism, Model Parallelism, Kernel Optimization |
Composer |
@@
Fully Sharded Data Parallelism, Elastic sharded checkpointing, Flash Attention |
vLLM |
@@
Data Parallelism, Model Parallelism, Tensor Paralellism, Efficient management via PagedAttention, Optimized CUDA kernels, Dynamic Batching, Quantization |
OpenLLM |
@@
[…]
Distributed Finetuning and Inference, Integration with BentoML, LangChain, and Transformers Agents, Prometheus Metrics, Token Streaming |
Ray LLM |
[…]
Distributed Inference, Integration with Alpa, Prompt Batching, Quantization, Prometheus Metrics |
MLC LLM |
[…]
Distributed Inference, Compiler Acceleration, Prompt Batching, Quantization |
Sax |
[…]
Distribute Inference, Serves PaxML, JAX, and PyTorch models, Slice Serving, Prometheus Metrics |
Mosec |
[…]
Distribute Inference, Dynamic Batching, Rust-based Task Coordinator, Prometheus Metrics |
LLM Foundry |
[…]
Distribute Inference, Dynamic Batching, Prompt Batching |
4. LLM frameworks
@@
DeepSpeed. Developed by Microsoft, DeepSpeed (Rasley et al., 2020) is an integrated framework for both training and deploying LLMs. It has been used to train large models like Megatron-Turing NLG 530B (Smith et al., 2022) (in a joint effort with Nvidia Megatron framework) and BLOOM (Scao et al., 2022). Within this framework, DeepSpeed-Inference is the foundational library. A pivotal feature of this module is ZeRO-Inference (Rajbhandari et al., 2020, 2021b), an optimization technique created to address GPU memory constraints for large model inference. ZeRO-Inference distributes model states across multiple GPUs and CPUs, providing an approach to managing the memory constraints of individual nodes. Another aspect of DeepSpeed-Inference is its deep fusion mechanism, which allows for the fusion of operations without the necessity for global synchronization by tiling computations across iteration space dimensions (Ren et al., 2021; Tang et al., 2021; Li et al., 2021a; Lu et al., 2022b). Building on this, the DeepSpeed Model Implementations for Inference (DeepSpeed MII) module provides strategies for the deployment and management of popular deep learning models. Emphasizing performance, flexibility, and cost-efficiency, DeepSpeed MII incorporates advanced optimization techniques to improve model inference (Rajbhandari et al., 2021b; Yao et al., 2022a; Wu et al., 2023a). Furthermore, the introduction of DeepSpeed-Chat (Yao et al., 2023a) adds chat support to the ecosystem. This module focuses on training chatbot models across different scales, integrating techniques from Reinforcement Learning from Human Feedback (RLHF) (Griffith et al., 2013) with the DeepSpeed training system. Notably, its integration of the ZeRO-Offload optimizer (Ren et al., 2021) facilitates training on both CPUs and GPUs, irrespective of their memory capacities.
@@
Megatron. Megatron (Shoeybi et al., 2019) constitutes Nvidia’s efforts to streamline training and deployment of LLMs such as GPT (Radford et al., 2019) and T5 (Raffel et al., 2020). It is the underlying framework used for Nvidia’s Megatron models (Shoeybi et al., 2019; Narayanan et al., 2021; Korthikanti et al., 2023). Megatron encompasses various specialized tools and frameworks for Nvidia GPUs. Central to Megatron’s design is the strategic decomposition of the model’s tensor operations, distributed across multiple GPUs, to optimize both processing speed and memory utilization, thus enhancing training throughput without compromising model fidelity (Shoeybi et al., 2019). Megatron also uses FasterTransformer (NVIDIA, 2023) for optimizing the inference process for large Transformer models. Furthermore, FasterTransformer is used for handling varying precision modes like FP16 and INT8, catering to diverse operational needs. The system also incorporates algorithms tailored to specific GPU architectures like Turing and Volta, emphasizing performance optimization (NVIDIA, 2023). Finally, Megatron uses TensorRT-LLM which provides developers with advanced tools and optimizations specifically tailored for LLMs, aiming to significantly reduce latency and enhance throughput for real-time applications. Notably, TensorRT-LLM integrates optimized kernels from FasterTransformer (Timonin et al., 2022) and employs tensor parallelism, facilitating efficient inference at scale across multiple GPUs and servers without necessitating developer intervention or model changes.
Alpa. Alpa (Zheng et al., 2022) is a library for training and serving large-scale neural networks. Alpa strategically addresses both inter- and intra-operator parallelism, aiming for a holistic enhancement in distributed deep learning performance. It has example implementations of GPT-2 (Radford et al., 2019), BLOOM (Scao et al., 2022), OPT (Zhang et al., 2022d), CodeGen (Nijkamp et al., 2022) among others. At the core of Alpa’s methodology is its automatic parallelization. By deploying an auto-tuning framework, Alpa dynamically identifies the optimal parallelism strategy tailored to specific deep learning models and hardware configurations. Furthermore, Alpa showcases an integrated design that combines both data and model parallelism (Zhuang et al., 2023b; Li et al., 2023h). By doing so, Alpa harnesses the collective benefits of these parallelism techniques, leading to optimized resource utilization and enhanced training throughput during serving.
ColossalAI. ColossalAI (Li et al., 2023b) is a framework tailored to address the challenges of large-scale distributed training (Wang et al., 2021). ColossalAI provides a unified solution that harmonizes scalability, efficiency, and versatility. It has implementations for LLaMA (Touvron et al., 2023b), GPT-3 (Brown et al., 2020), GPT-2 (Radford et al., 2019), BERT (Devlin et al., 2019), PaLM, OPT (Zhang et al., 2022d), ViT (Dosovitskiy et al., 2021). Central to Colossal-AI’s design is its emphasis on holistic integration. By amalgamating various components of deep learning pipelines, from data preprocessing to model training and validation, ColossalAI provides a streamlined platform that reduces fragmentation and enhances workflow efficiency (Bian et al., 2021). This integrated approach mitigates the complexities often associated with orchestrating large-scale training in distributed environments. Furthermore, recognizing the dynamic landscape of deep learning research and applications, the system is architected to be inherently modular (Chen et al., 2016). In addition, the framework integrates several other advanced optimization techniques (Bian et al., 2021; Li et al., 2021b; Wang et al., 2022f; Fang et al., 2022; Fang et al., 2023; Liu et al., 2023c) and features like quantization, gradient accumulation, and mixed precision. By leveraging state-of-the-art algorithms and methodologies, Colossal-AI seeks to optimize both computational and communication overheads inherent in parallel training, leading to reduced training times and enhanced model performance.
FairScale. Developed by Meta, FairScale (FairScale authors, 2021) is an extension library to PyTorch, dedicated to high-performance and large-scale training initiatives. The ethos of FairScale is rooted in three fundamental principles: usability, which emphasizes the ease of understanding and utilization of FairScale’s APIs aiming to minimize cognitive overhead for users; modularity, which endorses a seamless amalgamation of multiple FairScale APIs within the users’ training loops, thus promoting flexibility; and performance, which is centered around delivering optimal scaling and efficiency through FairScale’s APIs. Additionally, FairScale provides support for Fully Sharded Data Parallel (FSDP) as the preferred method for scaling the training operations of extensive neural networks. It is therefore a powerful tool for distributed training and inference. Additionally, it has key features for training in resource-constrained systems providing support for activation checkpointing, efficient model offloading, and scaling.
@@
Pax. Developed by Google, Pax (Authors, 2023a) is a JAX-based efficient distributed training framework. Pax has been used to train PaLM-2 (Anil et al., 2023) and Bard (Hsiao et al., 2023). It targets scalability and has reference examples for large model training, including across modalities (such as text, vision, speech, etc.). It is heavily integrated with JAX and uses many libraries in the JAX ecosystem. Pax contains many key components, including SeqIO to handle sequential data processing, Optax for optimization, Fiddle for configuration, Orbax for checkpointing, PyGLove for automatic differentiation, and Flax for creating high-performance neural networks.
@@
Composer. Designed by Mosaic ML, Composer (MosaicML, 2023a) is aimed at making the training of neural networks faster and more efficient. It has been used to train Mosaic ML’s MPT 7B and MPT 30B models and Replit’s Code V-1.5 3B. The library is built on top of PyTorch and provides a collection of speedup methods that users can incorporate into their own training loops or use with the Composer trainer for a better experience. It supports FSDP for efficient parallelism, elastic shared checkpointing for robust intermittent training, and a dataset streaming implementation allowing to download datasets from cloud blob storage on the fly during training. Composer is therefore designed to be versatile with a Functional API for integrating methods directly into its training loops, as well as a Trainer API which automatically implements a PyTorch-based training loop, reducing the workload for ML developers.
vLLM. vLLM (Kwon et al., 2023) represents a methodological shift in the approach to serving LLMs. Central to vLLM’s design is PagedAttention, a mechanism that segments the attention key and value (KV) cache for a set number of tokens. Unlike contiguous space storage, PagedAttention’s blocks for the KV cache are stored flexibly, akin to the virtual memory management. This facilitates memory sharing at a block level across various sequences tied to the same request or even different requests, thus enhancing memory management efficiency in handling attention mechanisms. It also allows on-demand buffer allocation, while also eliminating external fragmentation as the blocks are uniformly sized. Furthermore, vLLM incorporates an adaptive loading technique. This technique, rooted in heuristic methodologies, discerns the number of pages to be loaded into memory based on the input. Complementing this, vLLM integrates a parameter compression strategy as well. By storing model parameters in a compressed state and decompressing them during real-time serving, vLLM further optimizes memory usage. Additionally, vLLM supports state-of-the-art quantization techniques and optimized CUDA kernels supporting fast model execution. The library also added support for AMD’s ROCm GPUs. vLLM is therefore, not only a useful tool for distributed training, it can also handle efficient high-throughput model serving workloads.
OpenLLM. OpenLLM (Pham et al., 2023) delineates a comprehensive approach to the deployment and operation of LLMs within production environments. Anchored within the BentoML ecosystem, OpenLLM is crafted to bridge the gap between the training of LLMs and their seamless integration into real-world applications. A defining characteristic of OpenLLM is its emphasis on modularity and scalability. Recognizing the diverse needs of production environments, OpenLLM promotes a component-based architecture. Further enhancing its value proposition, OpenLLM integrates advanced caching mechanisms. By leveraging these mechanisms, the system aims to optimize repetitive queries, leading to reduced operational costs and enhanced response times. Additionally, OpenLLM’s design incorporates robust monitoring and logging tools, ensuring that operational insights are readily available for performance tuning and troubleshooting.
Ray-LLM. Ray-LLM (Project, 2023) represents a strategic fusion of LLMs with the Ray ecosystem (Moritz et al., 2018), aiming to optimize the deployment and operation of LLMs. Situated at the intersection of cutting-edge model architecture and scalable infrastructure, Ray-LLM seeks to redefine the paradigms of LLM utilization. At the core of Ray-LLM’s approach is the leveraging of Ray’s inherent distributed computing capabilities. Recognizing the computational demands of LLMs, Ray-LLM integrates Ray’s distributed task scheduling and execution mechanisms, ensuring that LLM tasks are efficiently distributed across available resources. This seamless integration potentially leads to enhanced model performance, reduced latency, and optimized resource utilization. Since it is built on top of the Ray Ecosystem, Ray-LLM is a good library to quickly prototype, train and deploy large models on clusters. It also comes with advanced monitoring support as well, enabling its usage in serving.
@@
MLC-LLM. MLC-LLM (team, 2023) aspires to empower individuals to develop, optimize, and deploy AI models on a diverse array of devices. Central to MLC-LLM’s approach is the concept of device-native AI. Recognizing the vast spectrum of devices in use today, from high-end servers to smartphones, MLC-LLM compiles models and deploys them in a process that is inherently tailored to the specific capabilities and constraints of each device (Chen et al., 2018; Shao et al., 2022; Feng et al., 2023). This device-native focus ensures that AI models are not only efficient but also highly optimized for the environments in which they operate. With its strong focus on compiling and optimizing models for prototyping on edge devices, MLC-LLM is a powerful tool for deploying on-device AI models and exhibits state-of-the-art performance in terms of throughput across a range of devices.
Sax. Sax (Authors, 2023b) is a platform designed by Google for deploying Pax, JAX, and PyTorch models for inference tasks. Within Sax, there is a unit referred to as Sax cell (or Sax cluster) thatl is made up of an administrative server coupled with multiple model servers. The role of the admin server is multifaceted: it monitors the model servers, allocates published models to these servers for inference, and guides clients in finding the appropriate model server for specific published models. Sax is essentially complementary to the Pax framework and while Pax focuses on massively distributed workloads, Sax is geared toward model serving.
@@
Mosec. Mosec (Yang et al., 2021) is designed for serving large deep learning models particularly in cloud environments. It is built to streamline the serving of machine learning models into backend services and microservices. Key features include high performance due to Rust-built web layer and task coordination, easy-to-use Python interface, dynamic batching, pipelined stages for handling mixed workloads, and cloud-friendliness with model warmup, graceful shutdown, and Prometheus monitoring metrics, making it easily manageable by Kubernetes or other container orchestration systems. Mosec is centered around cloud ecosystems and is well suited for serving models efficiently with its web layer, allowing developers to focus on model optimization and backend logic.
LLM Foundry. LLM Foundry (MosaicML, 2023b) is a library for finetuning, evaluating, and deploying LLMs for inference with Composer and the MosaicML platform. It supports distributed inference, dynamic batching, and prompt batching for efficient deployment. Similar to its complimentary training framework Composer, LLM Foundry is designed to be easy to use, efficient, and flexible, aimed at enabling rapid experimentation with the latest techniques in LLMs. It also provides straightforward interfaces to Mosaic’s Pretrained Transformers (MPT) (GPT-style models with built-in support for features like FlashAttention (Dao et al., 2022) and ALiBi (Press et al., 2022)). It is complementary to MosaicML’s Composer framework and while Composer focuses on distributed training, LLM Foundry provides support for deploying those models and enabling rapid experimentation with the latest techniques.
5. Concluding Remarks

### v4

@@ 3.2.2 Prompt Compression
Prompt compression (Figure 20(a)) accelerates the processing of LLM inputs through either condensing lengthy prompt inputs or learning compact prompt representations. Mu et al. (2023) propose to train LLMs to distill prompts into a more concise set of tokens, referred to as gist tokens. These gist tokens encapsulate the knowledge of the original prompt and can be stored for future use. In doing so, it is able to compress prompts by up to 26x, leading to a reduction in floating-point operations per second (FLOPs) by up to 40%. Chevalier et al. (2023) propose AutoCompressors to condense long textual contexts into compact vectors, known as summary vectors, which can then be used as soft prompts for the language model. These summary vectors extend the model’s context window, allowing it to handle longer documents with much less computational cost. In particular, AutoCompressors can utilize long contexts to improve perplexity of both fine-tuned OPT and LLaMA-2 on sequences of up to 30,720 tokens. Jung & Kim (2023) propose Prompt Compression with Reinforcement Learning (PCRL) that employs a policy network to directly edit prompts, aiming to reduce token count while preserving performance. It achieves an average reduction of 24.6% in token count across various instruction prompts. Ge et al. (2024) propose In-context Autoencoder (ICAE), which consists of a learnable encoder and a fixed decoder. The encoder compresses a long context into a limited number of memory slots, which the target language model can then condition on. With such design, ICAE is able to obtain 4x context compression. Nugget 2D (Qin et al., 2023b) represents the historical context as compact “nuggets” that are trained to enable reconstruction. Furthermore, it has the flexibility to be initialized using readily available models like LLaMA. Eperimental results show that Nugget 2D compresses context at a 20x compression ratio with a BLEU score of 98% for reconstruction, achieving nearly lossless encoding. Lastly, LongLLMLingua (Jiang et al., 2024) introduces a prompt compression technique containing question-aware coarse-to-fine compression, document reordering, dynamic compression ratios, and post-compression sub-sequence recovery to enhance LLMs’ key information perception. Experimental results show that LongLLMLingua achieves 17.1% better performance over the original prompt with 4x fewer tokens as input to GPT-3.5-Turbo.
3.2.3 Prompt Generation
@@
Prompt generation (Figure 20(b)) enhances efficiency by automatically creating effective prompts that guide the model in generating specific and relevant responses. For instance, AutoPrompt (Shin et al., 2020) proposes an automated method to generate prompts for a diverse set of tasks based on a gradient-guided search. It underscores the significance of human-written text in refining the quality and authenticity of data, emphasizing its pivotal role in optimizing LLM performance. Experimental results demonstrate that AutoPrompt outperforms manually created prompts on the LAMA benchmark in eliciting more precise factual knowledge from LLM. TempLM (Zhang et al., 2023e) proposes to combine generative and template-based methodologies to distill LLMs into template-based generators, offering a harmonized solution for data-to-text tasks. TempLM not only reduces the unfaithfulness rate of a fine-tuned BART model from 83% to 0%, but also substantially improves upon human-written ones in BERTScore. Lastly, PromptGen (Zhang et al., 2022c) considers dynamic prompt generation for knowledge probing based on pre-trained LLMs. It automatically generates prompts conditional on the input sentence and outperforms AutoPrompt on the LAMA benchmark.
4 LLM Frameworks
LLM frameworks can be in general grouped based on whether they support the tasks of training, fine-tuning, and inference. Specifically, frameworks that support training and/or fine-tuning aim to provide scalable, efficient, and flexible infrastructure that improves computation efficiency, reduces memory footprint, optimizes communication efficiency, and ensures reliability of the training/fine-tuning process. Frameworks that support inference focus on optimizing inference throughput and reducing memory footprint and latency. These frameworks offer a variety of deployment features to serve LLM requests. Table LABEL:tab:llmframeworks provides a summary of existing LLM frameworks along with their key features.
DeepSpeed. Developed by Microsoft, DeepSpeed (Rasley et al., 2020) is an integrated framework for both training and serving LLMs. It has been used to train large models like Megatron-Turing NLG 530B (Smith et al., 2022) (in a joint effort with Nvidia Megatron framework) and BLOOM (Scao et al., 2023). Within this framework, DeepSpeed-Inference is the foundational library. A pivotal feature of DeepSpeed-Inference is ZeRO-Inference (Rajbhandari et al., 2020; 2021), an optimization technique created to address GPU memory constraints for large model inference. ZeRO-Inference distributes model states across multiple GPUs and CPUs, providing an approach to managing the memory constraints of GPUs.
caption = Comparison of LLM frameworks.,
label = tab:llmframeworks,
width = colspec = Q[140]Q[90]Q[200]Q[90]Q[500],
columneven = c,
column3 = c,
cell15 = c,
hline1-2,19 = -,
Framework & Training Fine-Tuning Inference Key Features
@@ 4 LLM Frameworks
Another important feature of DeepSpeed-Inference is its deep fusion mechanism, which allows for the fusion of operations without the necessity for global synchronization by tiling computations across iteration space dimensions (Ren et al., 2021; Tang et al., 2021; Li et al., 2022; Lu et al., 2023). Building on this, the DeepSpeed Model Implementations for Inference (DeepSpeed MII) module introduces Dynamic SplitFuse (Holmes et al., 2024), which leverages continuous batching (Yu et al., 2022) and non-contiguous KV caches to enable increased occupancy and higher responsivity for LLM serving. It also supports scaling beyond tensor, data, and pipeline parallelism (3D Parallelism) with Zero Infinity (Rajbhandari et al., 2021) and efficient post-training quantization to Int8 with ZeroQuant (Yao et al., 2022b), Int4 (W4A4) with Wu et al. (2023a) and ternary quantization with XTC (Wu et al., 2022b). Furthermore, the introduction of DeepSpeed-Chat (Yao et al., 2023c) adds chat support to the framework. This module focuses on training chatbot models across different scales, integrating techniques like Reinforcement Learning from Human Feedback (RLHF) (Griffith et al., 2013) with the DeepSpeed training system. Notably, its integration of the ZeRO-Offload optimizer (Ren et al., 2021) facilitates training on both CPUs and GPUs, irrespective of their memory capacities.
@@
Megatron. Megatron (Shoeybi et al., 2020) is Nvidia’s efforts to streamline training and serving of LLMs such as GPT (Radford et al., 2019) and T5 (Raffel et al., 2020). It is the underlying framework used for Nvidia Megatron models (Shoeybi et al., 2020; Narayanan et al., 2021; Korthikanti et al., 2023). Central to Megatron’s design is the strategic decomposition of the model’s tensor operations, distributed across multiple GPUs, to optimize both processing speed and memory utilization, thus enhancing training throughput without compromising model quality (Shoeybi et al., 2020). In conjunction with 3D Parallelism, it also implements sequence parallelism and selective activation recomputation (Korthikanti et al., 2023), which enhances training efficiency. Megatron also uses FasterTransformer (NVIDIA, 2023a) for optimizing the inference process for large Transformer models and handling varying precision modes like FP16 and INT8, catering to diverse operational needs.
Colossal-AI. Colossal-AI (Li et al., 2023c) is a framework mainly designed for large-scale distributed training (Wang et al., 2023a). Colossal-AI unifies a wide range of parallelism techniques (Xu & You, 2023; Wang et al., 2023a; Bian et al., 2021) including sequence parallelism (Li et al., 2023d), auto-parallelism (Liu et al., 2023c), and Zero Redundancy Optimizer (Zhao et al., 2022). It also implements heterogeneous memory management (Fang et al., 2023) through a streamlined API. This integrated approach mitigates the steep learning curve often associated with orchestrating large-scale training in distributed environments. In addition, the framework integrates several other features like quantization (Frantar et al., 2023; Xiao et al., 2023), RLHF (Griffith et al., 2013), OpenMoE, and mixed-precision training.
Nanotron. Nanotron (HuggingFace, 2023a), introduced by Huggingface, is an LLM training framework with a primary focus on providing functionality with minimal overhead. As such, the library only subjectively incorporates the best optimizations required for modern LLM training requirements. Some highlights are its implementation of tensor, data, and pipeline parallelism with one-forward-one-backward pipeline engine, ZeRO-1 optimizer (Zhao et al., 2022), FP32 gradient accumulation, parameter tying/sharding and spectral µTransfer parametrization (Yang et al., 2022) for scaling up neural networks. Nanotron also incorporates DoReMi (Xie et al., 2023a) to further speed up training.
MegaBlocks. Developed by Databricks, MegaBlocks (Gale et al., 2023) is an LLM training framework for training Mixture-of-Experts (MoE) models. The core of MegaBlocks is dropless-MoE, a reformulation of MoE in terms of block-sparse operations, that avoids token dropping without sacrificing hardware efficiency. This design simplifies and accelerates training as it does not require the capacity factor as a hyper-parameter.
FairScale. Developed by Meta, FairScale (FairScale authors, 2021) is an extension library to PyTorch, dedicated to high-performance and large-scale training. As a highlight, FairScale uses Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023c) as the preferred method for scaling the training operations of large neural networks. It uses AdaScale optimizer (Johnson et al., 2020) as its distributed optimizer.
@@
Pax. Developed by Google, Pax (Google, 2023a) is a JAX-based distributed training framework. Pax has been used to train PaLM-2 (Anil et al., 2023) and Bard (Hsiao et al., 2023). It targets scalability and has reference examples for large model training, including across modalities (e.g., text, vision, speech). It supports data and tensor parallelism, and is heavily integrated with JAX and uses many libraries in the JAX ecosystem. Several key components Pax contains include SeqIO to handle sequential data processing, Optax for optimization, Fiddle for configuration, Orbax for checkpointing, PyGLove for automatic differentiation, and Flax for creating high-performance neural networks.
@@
Composer. Composer (MosaicML, 2023a) is an LLM framework designed by Mosaic ML. It has been used to train Mosaic ML’s MPT 7B and MPT 30B models and Replit’s Code V-1.5 3B. The framework is built on top of PyTorch and provides a collection of acceleration methods that users can incorporate into their training loops or use with the Composer trainer. Composer supports FSDP for efficient parallelism, elastic shared checkpointing for robust intermittent training, and a dataset streaming implementation allowing the download of datasets from cloud blob storage on the fly during training. Composer also provides a functional API for integrating methods directly into its training loops, as well as a Trainer API which automatically implements a PyTorch-based training loop, reducing the workload for ML developers.
OpenLLM. OpenLLM (Pham et al., 2023) was made by BentoML to serve and fine-tune LLMs within production environments. Anchored within the BentoML ecosystem, OpenLLM makes it easy to self-host LLMs and integrate them with other cloud services. OpenLLM emphasizes on flexibility, SOTA LLM support, and streamlined APIs for self-hosting. Recognizing the diverse needs of production environments, OpenLLM supports the automatic generation of docker images. It also supports serving models as serverless endpoints using BentoML’s cloud platform. OpenLLM further integrates advanced quantization techniques (GPTQ (Frantar et al., 2023), AWQ (Lin et al., 2023), SqueezeLLM (Kim et al., 2024), SpQR (Dettmers et al., 2024), LLM.int8 (Dettmers et al., 2022)) for efficient inference. OpenLLM’s design also incorporates robust monitoring with Prometheus metrics and logging tools, ensuring that operational insights are readily available for performance tuning and troubleshooting.
LLM Foundry. LLM Foundry (MosaicML, 2023b) is a library developed by MosaicML for fine-tuning, evaluating, and serving LLMs. It supports distributed inference via FSDP and continuous batching (Yu et al., 2022) for efficient serving. It also has cloud integration with the MosaicML platform.
vLLM. vLLM (Kwon et al., 2023) is an open-source library for LLM inference and serving. It adopts a different design in how KV cache is stored in memory. Central to this design is PagedAttention, a mechanism that segments the attention key and value (KV) cache for a set number of tokens. Unlike contiguous space storage, PagedAttention’s blocks for the KV cache are stored flexibly, similar to the virtual memory management in operating systems. This facilitates memory sharing at a block level across various sequences tied to the same request or even different requests, thus enhancing memory management efficiency in handling attention mechanisms. Hence, vLLM can reduce the total memory usage when using complex decoding techniques such as parallel sampling and beam search as the memory blocks can be shared across different candidate samples. It also allows on-demand buffer allocation, while also eliminating external fragmentation as the blocks are uniformly sized. Furthermore, vLLM incorporates safeguards that prevent GPU memory overflow due to the increasing size of KV cache by evicting and recovering blocks as needed via swapping and recomputation. vLLM also supports Multi-LoRA (Wang et al., 2023g), continuous batching (Yu et al., 2022) and quantization (GPTQ (Frantar et al., 2023), AWQ (Lin et al., 2023) and SqueezeLLM (Kim et al., 2024)). Lastly, it implements efficient kernels for both Nvidia and AMD GPUs.
TensorRT-LLM. TensorRT-LLM (NVIDIA, 2023b) is a streamlined library to serve LLMs. Built on top of the TensorRT engine, TensorRT-LLM integrates optimized kernels from FasterTransformer (Timonin et al., 2022) and employs tensor parallelism, facilitating efficient inference at scale across multiple GPUs and servers without necessitating developer intervention or model changes. It also integrates seamlessly with the Nvidia Triton Inference Server for serving LLMs. Additionally, it offers features like tensor parallelism, continuous batching (Yu et al., 2022), and PagedAttention (Kwon et al., 2023) as well as supports a wide range of quantization modes and techniques.
Text-Generation-Inference (TGI). Text-Generation-Inference (TGI) (HuggingFace, 2023b) is a high-performance LLM serving library developed by Huggingface and is used to power Hugging Chat. TGI supports a large variety of LLMs, and offers a wide range of features like tensor parallelism, continuous batching (Yu et al., 2022), efficient attention mechanisms like FlashAttention (Dao et al., 2022) and PagedAttention (Kwon et al., 2023), and quantization (BitsAndBytes (Dettmers, 2023), GPTQ (Frantar et al., 2023)) support.
RayLLM. RayLLM (Ray Project, 2023) is an LLM serving framework as part of the Ray ecosystem (Moritz et al., 2018). Built with Ray Serve and the Ray library developed by AnyScale, RayLLM eases LLM serving in multi-GPU and multi-node systems. At the core of RayLLM is the leveraging of Ray’s inherent distributed computing capabilities. RayLLM integrates Ray’s distributed task scheduling and execution mechanisms, ensuring that LLM tasks are efficiently distributed across available resources. RayLLM also simplifies adding new LLMs for custom use cases, and offers auto-scaling support and high-performance features like continuous batching (Yu et al., 2022), quantization (GPTQ (Frantar et al., 2023), AWQ (Lin et al., 2023), SqueezeLLM (Kim et al., 2024)), and monitoring via Prometheus metrics. It also comes with advanced monitoring support as well which includes a CLI and a web frontend that can be used to compare and rank models and get cost and latency estimates.
@@
MLC LLM. MLC LLM (Machine Learning Compilation for Large Language Models) (MLC-LLM, 2023) allows individuals to develop, optimize, and deploy LLMs on a wide range of platforms such as mobile phones and web browsers. Central to MLC LLM is its focus on machine learning compilation techniques. MLC-LLM compiles LLMs and deploys them in a process that is inherently tailored to the specific capabilities and constraints of each platform and hardware (Chen et al., 2018; Shao et al., 2022; Feng et al., 2023). This platform-native approach ensures that LLMs are not only efficient but also highly optimized for the platforms in which they operate.
Sax. Sax (Google, 2023b) is a platform designed by Google for serving Pax, JAX, and PyTorch models for inference tasks. It supports distributed inference with tensor and data parallelism. It also integrates easily with Google Cloud and cloud monitoring with Prometheus metrics.
@@
Mosec. Mosec (Yang et al., 2021) is a serving framework built to streamline the serving of machine learning models into backend services and microservices. Mosec’s key features include continuous batching (Yu et al., 2022), pipelined stages for handling mixed workloads, and essential cloud features like model warmup, graceful shutdown, and Prometheus monitoring metrics, making it easily manageable by Kubernetes or other container systems.
5 Concluding Remarks